Find all the points on the graph of (x^2 + y^2)^2 = x^2−y^2 that have horizontal tangent line.


1
Expert's answer
2021-05-04T11:40:16-0400

Horizontal tangent at the point "x=x_0" requires "y'(x=x_0)=0".

Let's apply operator "\frac{d}{dx}" to our expression, "y" being function "y(x)":

"x^4 +2x^2y^2+y^2=x^2-y^2 \quad \bigg | \frac{d}{dx}"

"4x^3+4xy^2+4x^2yy'+4y^3y'=2x-2yy'"

Substitute "x=x_0" into this expression, remembering "y'(x=x_0)=0" :

"4x_0^3+4x_0y(x_0)=2x_0"

Let "y(x_0)=y_0". We can also substitute "x=x_0" into our initial expression. Then we get the system of two equations and "y_0" and "x_0" can be found.

"\begin{cases}
4x_0^3+4x_0y_0^2=2x_0\\
(x_0^2+y_0^2)^2=x_0^2-y_0^2
\end{cases}\\
\begin{cases}
x_0^2+y_0^2=\frac{1}{2}\\
(x_0^2+y_0^2)^2=x_0^2-y_0^2
\end{cases}\\
\begin{cases}
x_0^2=\frac{1}{2}-y_0^2\\
\frac{1}{4}=\frac{1}{2}-2y_0^2
\end{cases}\\
\begin{cases}
x_0^2=\frac{3}{8}\\
y_0^2=\frac{1}{8}
\end{cases}"

So we got four points where tanget is horizontal: "(\frac{1}{\sqrt{8}},\sqrt{\frac{3}{8}}), (-\frac{1}{\sqrt{8}},\sqrt{\frac{3}{8}}), (\frac{1}{\sqrt{8}},-\sqrt{\frac{3}{8}}), (-\frac{1}{\sqrt{8}},-\sqrt{\frac{3}{8}})"